OpenStudy (anonymous):

payments will be 244.22 per month for 8 years the interest rate is 7.35% what is the asking price?

OpenStudy (anonymous):

Here is a simplified version of your question. You pay $50/month for 3 months with 10% interest rate. Total Payment is $50*3 = $150. 3rd month \[50 = X3 \times 1.1\]\[X3 = 50/1.1 = 45.45\] 2nd month \[45.45+50 = 95.45 = X2\times1.1\]\[X2 = 95.45\div1.1=86.77\] 1st month \[86.77+50 = 136.77 = X1\]\[X1 (AskingPrice) = 136.77\] 136.77 is the asking price. There should be $13.23 of interest to make $150.00. 45.45*0.1 + 86.77*0.1 = 4.55+8.68 = $13.23 .